# GNSS-inertial state initialization by distance residuals

### v1

@@ Abstract
Initializing the state of a sensorized platform can be challenging, as a limited set of initial measurements often carry limited information, leading to poor initial estimates that may converge to local minima during non-linear optimization. This paper proposes a novel GNSS-inertial initialization strategy that delays the use of global GNSS measurements until sufficient information is available to accurately estimate the transformation between the GNSS and inertial frames. Instead, the method initially relies on GNSS relative distance residuals. To determine the optimal moment for switching to global measurements, we introduce a criterion based on the evolution of the Hessian matrix’s singular values. Experiments on the EuRoC and GVINS datasets show that our approach consistently outperforms the naive strategy of using global GNSS data from the start, yielding more accurate and robust initializations.
@@ I Introduction
State initialization is a well-known challenge in sensor fusion, as evidenced by extensive research in the visual-inertial domain [1, 2, 3, 4]. Early sensor measurements often carry limited information, and certain states may even be unobservable. As a result, initial estimates can be highly inaccurate, leading to convergence toward local minima rather than the global optimum. To address this, dedicated initialization strategies are typically developed.
In this paper, we show for the first time that state initialization may also pose a challenge in GNSS-inertial systems. We introduce a novel two-stage initialization method that postpones the use of global position measurements until they become sufficiently informative. To determine the appropriate timing, we derive a general criterion that evaluates the informativeness of the available data for accurate state estimation. During the initial stage, the method relies solely on relative distances between GNSS measurements, effectively mitigating inertial preintegration drift. Experimental results show that our approach consistently outperforms the naive strategy of fusing all measurements from the outset.
The fusion of GNSS and inertial sensors becomes particularly valuable in scenarios where each sensor individually faces significant limitations—motivating the need for our state initialization method. Global Navigation Satellite Systems (GNSS), such as GPS, have long been fundamental for positioning and navigation across a wide range of applications. However, their performance can degrade in challenging environments like dense urban areas or locations with obstructed satellite visibility. In addition, positioning accuracy is affected by several error sources, including satellite clock drift, atmospheric delays, and receiver noise [5].
Inertial sensors, such as accelerometers and gyroscopes, offer high-frequency measurements and are immune to external signal disruptions. Yet, they suffer from cumulative drift, which leads to increasing errors over time. Furthermore, certain motion patterns—such as constant velocity or straight-line trajectories—can result in incomplete observability of the inertial states, limiting estimation accuracy over extended sequences.
Figure 1: Illustration of our GNSS-IMU initialization approach. a) Global residuals remain inactive while the system lacks informative measurements. b) At time $k^{*}$, the GNSS-inertial measurements provide sufficient information to meet our criterion, enabling the inclusion of global constraints for accurate alignment.
@@ II Related Work
II-A GNSS, inertial and additional modalities
A wide range of methods have been developed to enhance GNSS/INS integration by incorporating additional sensors or leveraging learning-based techniques. One early contribution was the use of sigma-point Kalman filters for integrating GPS and inertial data, allowing more accurate estimation under nonlinear conditions and improving convergence from poor initial estimates [6]. Later, hybrid systems began to emerge. For example, Yao et al. [7] combined a Kalman filter with a refined multi-layer perceptron to estimate pseudo-GPS positions when satellite signals are unavailable. This model relates inertial observations to GPS increments, making it effective in GPS-denied environments. To address motion model uncertainty, Hu et al. [8] proposed a strong tracking unscented Kalman filter (RSTUKF), capable of dynamically adjusting its estimation based on the presence of model error, thereby improving accuracy under both nominal and degraded conditions. GNSS errors in urban environments have also been a major concern. A map-aided adaptive fusion method introduced by Atia et al. [9] uses dead-reckoning and map-matching to identify and correct faulty GNSS readings by exploiting known road constraints. Incorporating self-learning into traditional filters, Shen et al. [10] presented the SL-SRCKF, a square-root cubature Kalman filter with an online adaptation mechanism that maintains consistent performance under varying observation conditions. Similarly, Geragersian et al. [11] employed a GRU-based network to leverage historical motion data and reduce drift when GNSS signals are lost. Beyond GNSS and inertial data, visual information has also proven valuable. Cremona et al. [12] fused GNSS, stereo images, and IMU data into a tightly-coupled ORB-SLAM3 pipeline for agricultural robotics, significantly improving pose accuracy over traditional visual-inertial systems. GVINS [13] extended this approach to general mobile robotics, achieving smooth and consistent state estimation by tightly coupling all three modalities. More recently, Liu et al. [14] introduced InGVIO, a tightly-coupled GNSS-visual-inertial odometry pipeline using invariant filtering, and Lee et al. [15] proposed a comparable system tailored for high-accuracy localization under varying operational conditions.
II-B GNSS-inertial localization
@@
In the absence of vision or contextual information, GNSS-IMU fusion remains crucial, particularly in cost-sensitive or resource-constrained applications. He et al. [16] developed an adaptive robust Kalman filter that suppresses outliers in GNSS signals, ensuring stable integration in scenarios with signal degradation or interference. Kim et al. [17] addressed computational efficiency by a compressed fusion strategy that embeds GNSS/INS fusion within a SLAM-like architecture, maintaining reliable estimates with reduced complexity. Zhang et al. [18] proposed the KGP algorithm, which enhances error-state updates in INS/GPS integration through improved modeling, resulting in a more accurate and stable positioning across dynamic navigation tasks.
@@
Unlike methods based on linear models and filtering, our approach formulates the problem as a nonlinear optimization, providing increased flexibility and accuracy in modeling system dynamics. In the early stages, the formulation includes relative distance residuals between consecutive GNSS measurements, which ensure geometric consistency without requiring global alignment. Once the observability criterion is satisfied, these relative residuals are deactivated and replaced by absolute position residuals, enabling direct alignment to the global frame. This switch is governed by an observability-based mechanism that ensures global constraints are introduced only when the estimation problem is sufficiently well-conditioned, thereby improving trajectory consistency and overall robustness.
@@ III Notation and background
Throughout this article, bold lower-case letters ($\mathbf{x}$) represent vectors and bold upper-case letters ($\bm{\Sigma}$) represent matrices. Scalars will be represented by light lower-case letters (${\alpha}$), functions by light upper-case letters ($J$). Poses belongs to groups of rigid motion ${SE}(3)$ and are represented as $\mathbf{T}_{WB}=[\mathbf{R}_{WB},\hskip 1.42262pt^{W}\mathbf{p}]\in\mathbb{R}^{4\mathrm{x}4}$ which maps a point from references frame $B$ to world references frame $W$.
Optimization updates will be done in $\mathfrak{so}(3)$ and $\mathfrak{se}(3)$, defined as the tangent spaces of ${SO}(3)$, and ${SE}(3)$ respectively. The hat operator $(\cdot)^{\wedge}$ maps a vector $\bm{\omega}\in\mathbb{R}^{3}$ to its corresponding skew-symmetric matrix in $\mathfrak{so}(3)$:
$\bm{\omega}^{\wedge}=\begin{bmatrix}0&-\omega_{3}&\omega_{2}\\ \omega_{3}&0&-\omega_{1}\\ -\omega_{2}&\omega_{1}&0\end{bmatrix}$
Its inverse is the vee operator, $(\cdot)^{\vee}$, such that $(\bm{\omega}^{\wedge})^{\vee}=\bm{\omega}$.
Elements of $\mathfrak{se}(3)$ can be written as $\bm{\xi}=[\mathbf{t},\bm{\omega}]^{\top}\in\mathbb{R}^{6}$. The exponential map projects tangent vectors back to the manifold. For ${SO}(3)$, this is given by Rodrigues’ formula:
$\mathrm{Exp}(\bm{\omega}^{\wedge})=\mathbf{I}+\frac{\sin\theta}{\theta}\bm{\omega}^{\wedge}+\frac{1-\cos\theta}{\theta^{2}}(\bm{\omega}^{\wedge})^{2},\quad\theta=\|\bm{\omega}\|$
The exponential map on ${SE}(3)$ is
$\mathrm{Exp}_{{SE3}}(\bm{\xi})=\begin{bmatrix}\mathrm{Exp}(\bm{\omega}^{\wedge})&\mathbf{V}\mathbf{t}\\ \mathbf{0}^{1\times 3}&1\end{bmatrix}$
with
$\mathbf{V}=\mathbf{I}+\frac{1-\cos\theta}{\theta^{2}}\bm{\omega}^{\wedge}+\frac{\theta-\sin\theta}{\theta^{3}}(\bm{\omega}^{\wedge})^{2}$
The logarithmic map is the inverse of the exponential map
$\mathrm{Log}_{{SE3}}(\mathbf{T})=\begin{bmatrix}\mathbf{V}^{-1}\mathbf{t}\\ \mathrm{Log}(\mathbf{R})^{\vee}\end{bmatrix},~{}~{}\mathrm{Log}(\mathbf{R})=\frac{\theta}{2\sin\theta}(\mathbf{R}-\mathbf{R}^{\top})$
where $\theta=\cos^{-1}\left(\frac{\mathrm{tr}(\mathbf{R})-1}{2}\right)$, and
$\mathbf{V}^{-1}=\mathbf{I}-\frac{1}{2}\bm{\omega}^{\wedge}+\gamma(\bm{\omega}^{\wedge})^{2}\quad\text{with}\quad\gamma=\frac{1-\frac{\theta}{2}\cot(\theta/2)}{\theta^{2}}$
Linearizing rotational residuals involving the logarithmic map requires the inverse right Jacobian of ${SO}(3)$, denoted $\mathbf{J}_{r}^{-1}(\bm{\phi})$. This term accounts for the manifold’s curvature and appears explicitly in the inertial residuals. Its closed-form expression is provided in [19].
These tools allow us to perform consistent optimization over manifolds using retraction-based updates [20].
@@ III-A IMU dynamics
We can use the relative motion increment
$\Delta\mathbf{v}_{ij}\doteq\hskip 1.42262pt\mathbf{R}_{i}^{\top}(\mathbf{v}_{j}-\mathbf{v}_{i}-\mathbf{g}\Delta t_{ij}),$
in order to obtain a gravity vector estimation $\mathbf{g}^{*}$ as follows
${\mathbf{g}}^{*}=\frac{\mathbf{v}_{j}-\mathbf{v}_{i}}{\Delta t_{ij}}-\frac{\mathbf{R}_{i}\Delta\mathbf{v}_{ij}}{\Delta t_{ij}}$ (3)
@@ III-B Noise propagation
The covariance matrix of the raw IMU measurements noise $\bm{{\Sigma}_{\eta}}\in\mathbb{S}^{6}_{+}$ [^1] By $\mathbb{S}^{n}_{+}=\{\mathbf{\Sigma}\in\mathbb{R}^{n\times n}\ |\ \mathbf{\Sigma}=\mathbf{\Sigma}^{\top},\mathbf{\Sigma}\succeq 0\}$ we denote the set of $n\times n$ symmetric positive semidefinite matrices. is composed by sub-matrices $\bm{{\Sigma}_{\omega}},\bm{{\Sigma}_{\textbf{a}}}\in\mathbb{S}^{3}_{+}$
@@
where $\textbf{0}_{3\mathrm{x}3}$ is a $33\mathrm{x}$ 3 matrix for which all its elements are equal to zero.
@@
${\bm{\Sigma}}_{ij}=\mathbf{A}_{j-1}{\bm{\Sigma}}_{ij-1}\mathbf{A}_{j-1}^{\top}+\mathbf{B}_{j-1}\bm{{\Sigma}_{\eta}}\mathbf{B}_{j-1}^{\top}$ (5)
with initial conditions ${\bm{\Sigma}}_{ii}=\textbf{0}_{9\times 9}$ and $\mathbf{A}_{j-1}\in\mathbb{R}^{9\times 9}$ , $\mathbf{B}_{j-1}\in\mathbb{R}^{9\times 6}$ defined as
$\mathbf{A}_{j-1}=\begin{bmatrix}\Delta\tilde{\mathbf{R}}_{j-1j}^{\top}&\textbf{0}_{3\mathrm{x}3}&\textbf{0}_{3\mathrm{x}3}\\ -\Delta\tilde{\mathbf{R}}_{ij-1}(\tilde{\mathbf{a}}_{j-1}-\mathbf{b}_{i}^{a})^{\wedge}\Delta t&\textbf{I}_{3}&\textbf{0}_{3\mathrm{x}3}\\ -\frac{1}{2}\Delta\tilde{\mathbf{R}}_{ij-1}(\tilde{\mathbf{a}}_{j-1}-\mathbf{b}_{i}^{a})^{\wedge}\Delta t^{2}&\textbf{I}_{3}\Delta t&\textbf{I}_{3}\end{bmatrix}$
$\mathbf{B}_{j-1}=\begin{bmatrix}\mathbf{J}_{r}^{j-1}\Delta t&\textbf{0}_{3\mathrm{x}3}\\ \textbf{0}_{3\mathrm{x}3}&\Delta\tilde{\mathbf{R}}_{ij-1}\Delta t\\ \textbf{0}_{3\mathrm{x}3}&\frac{1}{2}\Delta\tilde{\mathbf{R}}_{ij-1}\Delta t^{2}\end{bmatrix}$
where $\textbf{I}_{3}$ stands for the identity matrix of size $3$.
For the computation of the above matrices, the preintegrated expressions in [21] were used. The matrix ${\bm{\Sigma}}_{ij}\in\mathbb{S}^{9}_{+}$ is composed by nine sub-matrices of dimension 3 by 3 each
[…]
The above matrix will be important in the calculation of inertial residuals in the next section.
@@ III-C Gravity direction representation
$\hat{\mathbf{g}}_{1}\boxplus\bm{\delta}=\mathbf{R}_{\hat{\mathbf{g}}_{1}}\,\mathrm{Exp}_{\mathbb{S}^{2}}(\bm{\delta})$
$\hat{\mathbf{g}}_{2}\boxminus\hat{\mathbf{g}}_{1}=\mathrm{Log}_{\mathbb{S}^{2}}(\mathbf{R}_{\hat{\mathbf{g}}_{1}}^{\top}\hat{\mathbf{g}}_{2})$
Here, $\mathbf{R}_{\hat{\mathbf{g}}_{1}}\in\mathrm{SO}(3)$ is a rotation matrix that aligns $\hat{\mathbf{g}}_{1}$ with the vertical axis, ensuring that perturbations remain tangent to the manifold. The explicit derivations of these operators are provided in [23].
This parametrization enables a minimal and numerically stable representation of gravity direction perturbations while preserving the unit-norm constraint. Unit vectors $\hat{\mathbf{g}}_{i}$ can be locally parameterized using 2D perturbations in the tangent space at $\hat{\mathbf{g}}$. The mapping from a local perturbation $\bm{\delta}\in\mathbb{R}^{2}$ to a variation in the ambient space $\mathbb{R}^{3}$ is given by the Jacobian $\mathbf{J}_{\mathbb{S}^{2}}\in\mathbb{R}^{3\times 2}$. This Jacobian is constructed from an orthonormal basis $\mathbf{B}$ of the tangent space, such that $\mathbf{J}_{\mathbb{S}^{2}}=\mathbf{B}$, with
@@
This construction ensures that perturbations remain on the tangent plane and respect the unit-norm constraint, in line with standard manifold-based estimation strategies [24, 25]. The matrix $\mathbf{J}_{\mathbb{S}^{2}}$ will be used in the derivation of the gravity residual Jacobians in the next section.

### v2

# GNSS-Inertial State Initialization Using Inter-Epoch Baseline Residuals

@@ Abstract
Initializing the state of a sensorized platform can be challenging, as a limited set of measurements often provide low-informative constraints that are in addition highly non-linear. This may lead to poor initial estimates that may converge to local minima during subsequent non-linear optimization. We propose an adaptive GNSS–inertial initialization strategy that delays the incorporation of global GNSS constraints until they become sufficiently informative. In the initial stage, our method leverages inter-epoch baseline vector residuals between consecutive GNSS fixes to mitigate inertial drift. To determine when to activate global constraints, we introduce a general criterion based on the evolution of the Hessian matrix’s singular values, effectively quantifying system observability. Experiments on EuRoC, GVINS and MARS-LVIG datasets show that our approach consistently outperforms the naive strategy of fusing all measurements from the outset, yielding more accurate and robust initializations.
@@ I Introduction
State initialization is a well-known challenge in sensor fusion, as evidenced by extensive research in the visual-inertial domain [1, 2, 3]. Early sensor measurements often carry limited information, and certain states may even be unobservable. As a result, initial estimates can be highly inaccurate, leading to convergence toward local minima rather than the global optimum. Although this problem has been widely studied in the visual-inertial domain, analogous challenges may arise in other sensor fusion modalities.
In this paper, we demonstrate that state initialization is equally critical in GNSS–inertial systems, where naive strategies may cause inconsistent estimates. We propose a two-stage initialization method that postpones the use of global position measurements until they become sufficiently informative. To determine the appropriate timing, we derive a general criterion that evaluates the informativeness of the available data for accurate state estimation. During the initial stage, the method relies solely on inter-epoch baseline vectors between GNSS measurements, effectively mitigating inertial preintegration drift. Experimental results show that our approach consistently outperforms the naive strategy of fusing all measurements from the outset. Note that we do not target steady-state navigation [4], but producing robust and accurate initialization estimates to reliably bootstrap UAVs and other mobile robots in field conditions.
The complementary nature of GNSS and inertial sensors motivates the need for such an initialization strategy. GNSS provides global positioning capabilities fundamental to modern navigation, but its accuracy can degrade under multipath, atmospheric effects, or obstructed satellite visibility [5]. Inertial sensors, by contrast, offer high-frequency motion cues that are immune to external disturbances, but their integration inevitably leads to cumulative drift. Moreover, under certain motion patterns—such as constant velocity or straight-line trajectories—inertial states may become only partially observable. These limitations underscore the importance of principled initialization methods that can exploit early measurements effectively and ensure robust fusion performance.
Figure 1: Illustration of our GNSS-inertial state initialization. a) Global residuals remain inactive while the system lacks informative measurements. b) At time $k^{*}$, the GNSS-inertial measurements provide sufficient information to meet our criterion, enabling the inclusion of global constraints for accurate alignment.
@@ II Related Work
II-A GNSS-inertial alignment
Several methods have been proposed for GNSS-inertial alignment, particularly in scenarios involving low-cost MEMS sensors and challenging motion conditions. Shin and El-Sheimy [6] introduced an Unscented Kalman filter (UKF) framework for in-motion alignment, capable of handling large initial attitude uncertainties without stationary initialization. Zhang et al. [7] later proposed a velocity-based optimization-based alignment (VBOBA) method that exploits GNSS velocity vectors to improve heading observability under low-dynamic conditions. Groves [8] provides a comprehensive treatment of transfer alignment techniques, where a high-grade reference system supports the initialization of a secondary INS in motion. Wu and Pan [9] further derived the velocity/position integration formula, laying theoretical foundations for subsequent optimization-based alignment techniques. These approaches remain highly relevant and complementary. In contrast, our method is designed to guarantee robust initialization that prevents divergence, rather than aiming at fine-grained in-motion alignment, and thus provides a different but compatible contribution to the GNSS-inertial initialization literature. Wu et al. [10] introduced RA-JAPE, a time-recursive Newton–Lagrange optimization that jointly estimates attitude, IMU biases, and the GNSS lever arm, enabling self-initialization and outperforming state-of-the-art methods.
II-B GNSS–IMU localization beyond filtering
@@
In the absence of vision or contextual information, GNSS-inertial fusion remains crucial, particularly in cost-sensitive or resource-constrained applications. He et al. [11] developed an adaptive robust Kalman filter that suppresses outliers in GNSS signals, ensuring stable integration in scenarios with signal degradation or interference. Kim et al. [12] addressed computational efficiency by a compressed fusion strategy that embeds GNSS-inertial fusion within a SLAM-like architecture, maintaining reliable estimates with reduced complexity. Zhang et al. [13] proposed the KGP algorithm, which enhances error-state updates in INS/GNSS integration through improved modeling, resulting in a more accurate and stable positioning across dynamic navigation tasks.
Beyond linear error-state filtering, factor-graph optimization (FGO) offers a flexible backend to fuse IMU preintegration with GNSS observations (code, carrier, Doppler), often yielding higher accuracy and resilience under degraded satellite geometry [14, 15]. Robustness to outliers and mis-modeling can be handled with robust costs and switchable constraints, which softly activate/deactivate factors based on consistency—a mechanism closely related to observability-aware gating [16].
@@
Unlike methods based on linear models and filtering, our approach formulates the problem as a nonlinear optimization, providing increased flexibility and accuracy in modeling system dynamics. In the early stages, the formulation includes inter-epoch baseline vectors (i.e., consecutive GNSS position differences) which ensure geometric consistency without requiring global alignment. Once full observability is achieved, these relative residuals are deactivated and replaced by absolute position residuals, enabling direct alignment to the global frame. This switch is governed by an observability-based mechanism that ensures global constraints are introduced only when the estimation problem is sufficiently well-conditioned, thereby improving trajectory consistency and overall robustness.
@@ III Notation and background
Throughout this article, bold lower-case letters ($\mathbf{x}$) represent vectors and bold upper-case letters ($\bm{\Sigma}$) represent matrices. Scalars will be denoted as light lower-case letters ($\alpha$}) and functions by light upper-case letters ($J$). Six-degrees-of-freedom motion will be represented by the rotation matrix $\mathbf{R}_{WB}\in SO(3)$ and translation vector $^{W}\mathbf{p}\in\mathbb{R}^{3}$.
Optimization is carried out on the tangent spaces $\mathfrak{so}(3)$ and $\mathfrak{se}(3)$ of $SO(3)$ and $SE(3)$, respectively, using the exponential and logarithmic maps for retraction-based updates [17]. We use the hat operator $(\cdot)^{\wedge}:\mathbb{R}^{3}\!\to\!\mathfrak{so}(3)$, which maps a 3D vector to its skew-symmetric matrix, and the vee operator $(\cdot)^{\vee}:\mathfrak{so}(3)\!\to\!\mathbb{R}^{3}$, its inverse. Closed-form expressions for the exponential/logarithmic maps and the inverse right Jacobian of $SO(3)$, required for linearizing rotational residuals, are detailed in [14].
@@ III-A IMU dynamics
Using relative motion increment $\Delta\mathbf{v}_{ij}\doteq\hskip 1.42262pt\mathbf{R}_{i}^{\top}(\mathbf{v}_{j}-\mathbf{v}_{i}-\mathbf{g}\Delta t_{ij}),$ a gravity estimation can be obtained
${\mathbf{g}}^{*}=\frac{\mathbf{v}_{j}-\mathbf{v}_{i}}{\Delta t_{ij}}-\frac{\mathbf{R}_{i}\Delta\mathbf{v}_{ij}}{\Delta t_{ij}}.$ (3)
@@ III-B Noise propagation
The covariance matrix of the raw IMU measurements noise $\bm{{\Sigma}_{\eta}}\in\mathbb{S}^{6}_{+}$ is composed [^1] By $\mathbb{S}^{n}_{+}=\{\mathbf{\Sigma}\in\mathbb{R}^{n\times n}\ |\ \mathbf{\Sigma}=\mathbf{\Sigma}^{\top},\mathbf{\Sigma}\succeq 0\}$ we denote the set of $n\times n$ symmetric positive semidefinite matrices. by sub-matrices $\bm{{\Sigma}_{\omega}},\bm{{\Sigma}_{\textbf{a}}}\in\mathbb{S}^{3}_{+}$
@@
where $\textbf{0}_{3\mathrm{x}3}$ is a $3\mathrm{x}3$ matrix for which all its elements are equal to zero.
@@
${\bm{\Sigma}}_{ij}=\mathbf{A}_{j-1}{\bm{\Sigma}}_{ij-1}\mathbf{A}_{j-1}^{\top}+\mathbf{B}_{j-1}\bm{{\Sigma}_{\eta}}\mathbf{B}_{j-1}^{\top},$ (5)
with initial conditions ${\bm{\Sigma}}_{ii}=\textbf{0}_{9\times 9}$. Matrices $\mathbf{A}_{j-1}\in\mathbb{R}^{9\times 9}$ and $\mathbf{B}_{j-1}\in\mathbb{R}^{9\times 6}$, together with the corresponding covariance propagation rules, are expressed explicitly in [18]. The noise matrix ${\bm{\Sigma}}_{ij}\in\mathbb{S}^{9}_{+}$ captures the uncertainty of rotation, velocity, and position increments and is a key component in defining the residuals in the next section.
@@ III-C Gravity direction representation
$\hat{\mathbf{g}}_{1}\boxplus\bm{\delta}=\mathbf{R}_{\hat{\mathbf{g}}_{1}}\,\mathrm{Exp}_{\mathbb{S}^{2}}(\bm{\delta}),\quad\hat{\mathbf{g}}_{2}\boxminus\hat{\mathbf{g}}_{1}=\mathrm{Log}_{\mathbb{S}^{2}}(\mathbf{R}_{\hat{\mathbf{g}}_{1}}^{\top}\hat{\mathbf{g}}_{2}),$
where $\mathbf{R}_{\hat{\mathbf{g}}_{1}}\in\mathrm{SO}(3)$ is a rotation matrix that aligns $\hat{\mathbf{g}}_{1}$ with the vertical axis, ensuring that perturbations remain tangent to the manifold. The explicit expressions are provided in [20]. This parametrization enables a minimal and numerically stable representation of gravity direction perturbations while preserving the unit-norm constraint. Unit vectors $\hat{\mathbf{g}}_{i}$ can be locally parameterized using 2D perturbations in the tangent space at $\hat{\mathbf{g}}$. The mapping from a local perturbation $\bm{\delta}\in\mathbb{R}^{2}$ to a variation in the ambient space $\mathbb{R}^{3}$ is given by the Jacobian $\mathbf{J}_{\mathbb{S}^{2}}\in\mathbb{R}^{3\times 2}$. This Jacobian is constructed from an orthonormal basis $\mathbf{B}$ of the tangent space, such that $\mathbf{J}_{\mathbb{S}^{2}}=\mathbf{B}$, with
@@
This construction ensures that perturbations remain on the tangent plane and respect the unit-norm constraint, in line with standard manifold-based estimation strategies [21, 22].